OpenStudy (anonymous):

Identify which of the following you should do first when evaluating the expression A. Multiply 33 and 2 B. Add 2 and 5 C. Take 5 to the third power D. Multiply 33 and x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (texaschic101):

where is the expression ?

OpenStudy (anonymous):

\[33x2+5^{2}\]

OpenStudy (nincompoop):

for A, it says to multiply 33 and 2 this will give us 33 x 2 = _______ for B, add 2 and 5 this will give us 2 + 5 = ________ for C, take 5 to the third power this will give us: \[5^3\] this mean 5 x 5 x 5 = _______ for D, multiply 33 and x this will give us\[33 \times x =\]

OpenStudy (anonymous):

A!

OpenStudy (nincompoop):

you didn't have an expression to solve so I took the liberty to show everything from A through D

OpenStudy (nincompoop):

there is what we call ORDER OF OPERATIONS this means that in an expression, we need to perform certain operations first before others. The simple order of operation has an acronym of PEMDAS, meaning we solve first what is inside the P for parenthesis, then followed by expressions with exponents (with powers or roots), M for multiplication and D for division, A for addition then S for subtraction. Note however that Multiplication and Division come in the order from left-to-right. If an expression has multiplication first, then we multiply before we divide. But if division came first, then we divide before multiply. The same goes with addition and subtraction.

OpenStudy (anonymous):

I was about to say 5^3 should be done first but wasn't sure between that and multiplying 33 * x

OpenStudy (anonymous):

Yeah I think he wrote it wrong it the 5^3 should be done first

OpenStudy (nincompoop):

your 32x2+5^2 does not look right nor the solution appears to be in the choices given. you might be mistaking the variable x with the multiplication sign. and the choice that involved exponent indicated that 5 be taken to the third power.